OpenStudy (anonymous):

You're going to use your TI-83 graphing calculator to compute the area between values of 7 and 9 for a distribution whose mean is 8 and whose standard deviation is 2. You would enter: A. normalcdf(7,9,8,2) B. normalpdf(7,9,8,2) C. invNorm(8,7,9) D. normalcdf(8,2,7,9) E. normalpdf(8,2,7,9)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

think back to the template normalcdf(a,b,mu,sigma) a = lower bound b = upper bound mu = mean sigma = standard deviation

jimthompson5910 (jim_thompson5910):

the first two numbers (a & b) are the bounds the next two are the mean and std dev

jimthompson5910 (jim_thompson5910):

if you leave out the mean and std dev, it assumes you're dealing with the standard normal distribution. So it uses the default values of mu = 0 and sigma = 1

jimthompson5910 (jim_thompson5910):

so which one is it

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

its either a or b

jimthompson5910 (jim_thompson5910):

do you see the difference in the two?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i dont know exactly how they r different

jimthompson5910 (jim_thompson5910):

the pdf is for single points on the curve the cdf is for areas under the curve

jimthompson5910 (jim_thompson5910):

you use cdf more I think

OpenStudy (anonymous):

oh so cdf is area

OpenStudy (anonymous):

so it is a

OpenStudy (anonymous):

ok cool thnxs

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

can i ask the next one on this thread ?

OpenStudy (anonymous):

A normal probability plot is used to: A. determine the area between two normal curve values. B. test a distribution for normalcy. C. help find the z-score corresponding to an area under a normal curve. D. convert raw data to standardized scores. E. compare raw data to z-scores.

jimthompson5910 (jim_thompson5910):

its fine either way

jimthompson5910 (jim_thompson5910):

there may be other uses for a probability plot, but the main use is to see if you're dealing with a normal distribution http://www.itl.nist.gov/div898/handbook/eda/section3/normprpl.htm if the transformed points all line on or close to a straight line, then the distribution is considered to be normal. Otherwise, it is not a normal distribution and it may be distributed in some other way.

OpenStudy (anonymous):

so its B

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

cool ty one more?

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

Consider normally distributed data with a mean of 35 and a standard deviation of 2. If there are 350 values between 31 and 33, how many values are there in the distribution? A. 2525 B. 2550 C. 2575 D. 2600 E. 2625

jimthompson5910 (jim_thompson5910):

first calculate the area under the curve from 31 to 33

OpenStudy (anonymous):

oh ok ithink i know how to do that

jimthompson5910 (jim_thompson5910):

yep use normalcdf

OpenStudy (anonymous):

i got 0???

OpenStudy (anonymous):

nvm its .135

jimthompson5910 (jim_thompson5910):

normalcdf(31,33,35,2) is what you should type in

jimthompson5910 (jim_thompson5910):

yes, so 0.135*x = 350, solve for x

OpenStudy (anonymous):

2592

jimthompson5910 (jim_thompson5910):

I'm getting the same

OpenStudy (anonymous):

so will it be 2600?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

I'm getting 0.135905122 for the result on the normal cdf, so, 0.135905122*x = 350 x = 350/0.135905122 x = 2,575.32604253134

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ty SOOOO much, will u be here in the future?

jimthompson5910 (jim_thompson5910):

yeah I should be